# Floralens: A Deep Learning Model for the Portuguese Native Flora

### v2

@@ Abstract
Keywords: automatic identification, citizen science, deep learning, computer vision
@@ 1 Introduction
In project Biolens [9], we are interested in creating lightweight identification models for web and mobile apps, possibly working offline (e.g., in the field) and having the option of not immediately sharing data (e.g., for research projects). Citizen science platforms do not provide this level of flexibility. Moreover, a detailed description of their methodological approach is lacking in the literature (cf. Section 2) making it difficult to use this accumulated experience for new projects. We therefore decided to independently develop a streamlined methodology to build datasets for native Portuguese species covering different taxa and to derive accurate CNN-based models using off-the-shelf machine-learning tools. Currently, we have four models: Lepilens and Mothlens (for butterflies and moths, together covering the order Lepidoptera); Dragonlens (for dragonflies and damselflies, covering the order Odonata), and, the latest addition, Floralens (described in this paper, covering the Plantae kingdom).
Figure 1: Detail of the FloraOn web application.
@@
The methodology used to derive all the Biolens models has two core traits: (a) the use of data from research-grade public repositories for dataset construction, and; (b) the use of Google’s AutoML Vision (GAMLV) to derive the actual models. This methodology is briefly described in a short scientific outreach article (in Portuguese) [12], and, more thoroughly, in MSc theses [13, 14] (both covering different stages of the work on Floralens), and a BSc project report [15] (covering Lepilens). Our evaluation of Floralens shows that, with a carefully designed dataset, current off-the-shelf machine-learning cloud-based services output models whose performance rivals that obtained with models provided by state-of-the-art citizen science projects.
The main contributions of this work are as follows:
  • a dataset for the Portuguese Flora based on published research-grade datasets, available from Zenodo [16];
@@
  • a high-accuracy GAMLV-based model for the Portuguese native flora publicly available via web and mobile applications; and
@@
  • a quantitative evaluation of the derived model and a comparison of its accuracy relative to the state-of-the-art platform Pl@ntNet.
@@
The remainder of this paper is structured as follows. Section 2 describes the current state-of-the-art regarding automatic taxonomic identification based on deep learning. Section 3 describes the construction of the dataset used in this study. Section 4 describes the generation of the model from the dataset using GAMLV. Section 5 describes the results obtained with the model. Section 6 describes the software artifacts and datasets produced in the scope of this work. Finally, Section 7 summarizes the main findings of this study and puts forward some future research goals.
2 Related Work
Convolutional Neural Networks (CNN) are deep neural networks composed of multiple layers of trainable convolutional nodes whose aggregate outputs are eventually fed to a final fully connected layer. They are often used for image classification. In this context, a convolution is an operation that applies a matrix known as a kernel to a given input matrix (an image or part of it). The kernel slides over the input, multiplying the overlapping matrix positions at each value and then summing the values. Depending on the form of the kernel, the resulting matrix can encode features such as edges, textures, and shapes, extracted from the original image. Recently, Vision Transformers emerged as powerful tools to derive accurate models for image classification tasks [17, 18, 19].
@@ 2 Related Work
The advent of deep learning allowed the development of the first tools for automatically identifying plant species from input images [20, 21, 22]. The success of these first efforts and their further refinement quickly reached a point in which automatic species identification rivaled identifications made by specialists [23], thus attesting to the transformative role of AI in this field [24]. Nowadays, models based on deep learning are central tools in major citizen science platforms such as iNaturalist [2], Observation.org [3] and Pl@ntNet [6] and are integrated in web and mobile applications. However, the training process and machine learning methodologies used by these citizen science platforms are not documented in detail in the literature: [2] merely makes reference to machine learning from an end-user perspective, some details can only be found online in scattered form [25]; [6] provides only a summary of the Pl@ntNet methodology to derive models, the information is also given in slightly more extended form in [26]; [3] is a short abstract concerning their mobile app ObsIdentify [27], and both the app and the site interface are enabled by an API [24, 28] whose underlying model is not explained.
@@
Deriving a CNN-based model or other ML models often requires expert knowledge, non-trivial configuration aspects, writing specialized code, and a heavyweight computational infrastructure for storing data and model training. The strain is more acute when the amount of data involved is non-negligible, as in the case of Floralens where we use approximately 300,000 photos to derive a model. AutoML cloud services address these needs, of which GAMLV is an example. Like Google, other major cloud providers have MLaaS offerings, e.g., Amazon Rekognition [29], Apple Create ML [30], and Azure AutoML [31]. These services allow the automation of several aspects of model derivation and are backed by computational infrastructures that include special-purpose hardware such as GPUs or TPUs, necessary to derive a model in a reasonable amount of time. Models derived using these tools are referenced in the literature of several areas of knowledge, e.g., biology [32], medicine [33, 34], agriculture [35], engineering [36]. Comparative results between GAMLV and one or more competitor platforms are provided in some of these works [32, 33, 34], generally illustrating good performance by GAMLV-derived models. We should note that our choice of GAMLV was solely motivated by the availability of Google Cloud research credits, which we could not obtain from other providers.
Citizen science platforms also generate important by-products in the form of their curated datasets, often made available to the public through biodiversity data portals, notably GBIF [11]. These datasets enable the development of other ML models as is the case of Floralens that, in addition to FloraOn, uses data sampled from GBIF datasets provided by iNaturalist, Observation.org, and Pl@ntNet (cf. Section 3).
Compared to the previously mentioned citizen science platforms whose domain of application is the global flora, Floralens is more specialized covering only the Portuguese native species. It is also not supported by a citizen science platform nor, for the time being, supports directly exporting data to them. Users can only export their observations into common data formats such as CSV and ZIP for images. In this respect, Floralens is closer in philosophy and implementation to the Flora Incognita Project [5]. The projects share the goal of devising machine learning models and related applications for flora identification while falling short of integration with citizen science platforms (e.g., iNaturalist, Pl@ntNet, Observation.org) or enabling data aggregation (e.g., GBIF).
To improve the identification precision, some platforms such as Pl@ntNet developed regional models by dividing the globe into several biogeographic domains [37] following published regional floras such as the WCVP/Kew [38], and introduced metadata providing information on the anatomic part of the plant depicted in an image, e.g., flower, leaf, stem. This extra metadata critically improves the precision of the models [39, 40]. At this stage, Floralens does not use image metadata to aid in the identification. A different trend that has recently generated considerable interest is the generation of global models based on extreme datasets, consisting of millions of images representing tens of thousands of individual species (for comparison, it is estimated that the world has $\sim$300K plant species), as exemplified by iNat Challenge [41] and PlantCLEF/LifeCLEF [42, 43].
3 Dataset Construction
We now describe the creation of the Floralens dataset, subsequently used for training the deep learning model.
3.1 Data Sources
@@
The FloraOn repository contains geo-referenced records of Flora species with associated images. The image data is relatively broad in scope, as it covers 78% of the entire catalog (2,127 out of 2,712 species), but has a limited volume: on average there are just 11 images per species, and, unsurprisingly, our 50-image lower bound threshold is not met for a single species. Hence, to adequately populate the Floralens dataset, we retrieved the FloraOn images but also considered image data from three publicly-available datasets stored and made available at GBIF by three citizen science platforms: iNaturalist [44], Observation.org [45], and Pl@ntNet [46].
Figure 2: The geographical region of interest for GBIF portal queries.
The GBIF datasets provide validated observation data and associated images, originally submitted by users of the respective platforms. However, the validation process differs among data sources, as discussed further in this section. For each dataset and each species in our universe, we used GBIF portal queries to obtain observation records in the Darwin Core Archive format [47]. Each such record corresponds to an observation of a specimen, typically made in the wild by citizen scientists or experts, accompanied by its taxonomic identification, its geographical location, and one or more images. The GBIF portal queries were parameterized to cover the European continent (Figure 2) as, in a preliminary analysis, we found that occurrence data from just Portugal or even the entire Iberian Peninsula would yield limited data in terms of volume and variety. Fortunately, many species of the Portuguese flora are widely distributed.
@@ 3.2 Assembly
Figure 3: Dataset histograms ($x$-axis: number of images; $y$-axis: number of species).
This source-based prioritization intends to define a dataset where images are less prone to identification errors. It takes into account the curation processes associated with each data source. FloraOn is curated by botanists and features high-quality images. These often feature subtle details that help secure the identification of a species. Pl@ntNet data goes through a curation process that involves machine learning, contributors’ reputation scores, and geo-based species verification [46]. Observation.org data can result from automatic validation through image recognition coupled with a check for other approved observations in the geographical vicinity, or through an expert volunteer when automated validation fails [45, 48]. Finally, iNaturalist identifications result from a crowd-sourcing effort whereby a “research-grade” identification for a photo can be obtained from the consensus of just two citizen scientists [44, 49].
We used photos without distinction of plant features as, in general, this metadata was not available from the curated datasets. Similarly, we do not attempt to gauge “photo quality”. The citizen-science platforms provide guidelines for submitting acceptable quality photos, and their curation processes mostly filter inadequate items (e.g., out of focus, under/over-exposure, presence of significant digital noise, insufficient resolution, subject not sufficiently separated from surrounding vegetation).
@@ 4 Model Derivation
Figure 4: Model derivation using GAMLV.
4.1 Preliminaries
@@
Figure 5: GAMLV interface for model training and deployment.
4.2 Training
@@ 4.3 Deployment
Figure 6: Layers of the CNN model (fragment).
4.4 Derived Model
@@
The CNN architectures at stake are picked from the MnasNet family [55], developed with mobile and embedded devices in mind. The high-level choice between models offered by GAMLV (back in Figure 5(a)) corresponds to three different MnasNet instantiations that do not differ in structure, just in the density of connections between layers and number of internal weights.
5 Results
We evaluated the Floralens model using the test split described in the previous section, hereafter designated by FLTS (Floralens test split), using standard metrics. We then complemented these baseline results with those obtained using two additional test sets described further in this section [56, 43]. Next, we briefly evaluated the model’s capability of identifying the genus (as opposed to the species), the rationale being that sometimes even if the species cannot be identified with good confidence, knowledge of the genus will be useful to the user. Finally, we compared the Floralens results for all test sets with Pl@ntnet models accessible through the Pl@ntNet API [57].
5.1 Method
The performance of the Floralens model was evaluated over several test datasets using the following metrics. Precision is the ratio of true positives (TP) relative to the total number of positives (TP $+$ FP). A positive (identification) occurs when the classification score (a probability) returned by the model equals or exceeds a confidence level set as the threshold for analysis. Recall is the ratio of true positives relative to the total number of true examples (TP $+$ FN). The confidence level is relevant for practical uses of a model. Using a low confidence level means that false positives are tolerable for the intended use, and we will get lower precision. On the other hand, a too high value for the confidence level will result in more false negatives hence lower recall.
@@ 5.1 Method
5.2 Baseline Results
In this and the following subsections, larger values indicate better model performance.
@@ 5.2 Baseline Results
Figure 7: FLTS results: precision and recall results.
Figure 7 shows the results for precision and recall for the Floralens model applied to the test set given in Table 2, more precisely the macro-average of precision and recall values for all species. The area-under-curve (AUC) for the precision-recall correlation (in 7(a)), also known as the average precision, is 0.72 (the maximum value would be 1.0). Putting the confidence levels in perspective (in 7(b)) we can visualize that precision and recall are approximately equal to 0.7 for a confidence level of 0.2. For a confidence level of 0.5 precision equals 0.85 and recall equals 0.53. Overall, the results indicate a reasonable predictive power for the Floralens model.
Table 3: FLTS results: Top-1, Top-5 and MRR.
Data source Top-1 Top-5 MRR
FloraOn 0.70 0.88 0.77
iNaturalist 0.70 0.87 0.77
Observation.org 0.64 0.83 0.72
Pl@ntNet 0.66 0.87 0.74
Overall 0.67 0.86 0.75
Table 3 lists the Top-1, Top-5, and MRR results for the FLTS, per data source used in the construction of the data set (Floralens, iNaturalist, Observation.org, and Pl@ntNet) and also in overall terms (last line in the table). The results indicate relatively homogeneous predictive power across all data sources, as the maximum difference in values between them does not exceed 0.06: a Top-1 value of 0.64 for Observation.org vs. corresponding values of 0.7 for FloraOn and iNaturalist. The overall measures again indicate reasonably good predictive power: 0.67 for Top-1 (i.e., roughly two-thirds of the FLTS images are correctly classified with rank 1), 0.85 for Top-5, and 0.75 for MRR.

### v3

@@ 1 Introduction
In project Biolens [9], we are interested in creating lightweight identification models for web and mobile apps, possibly working offline (e.g., in the field) and having the option of not immediately sharing data (e.g., for research projects). Most citizen science platforms do not provide this level of flexibility. Moreover, a detailed description of their methodological approach is lacking in the literature (cf. Section 2), making it difficult to use this accumulated experience for new projects. We, therefore, decided to independently develop a streamlined methodology to build datasets for native Portuguese species covering different taxa and to derive accurate CNN-based models using off-the-shelf machine-learning tools. Currently, Biolens includes four models: Lepilens and Mothlens (for butterflies and moths, together covering the order Lepidoptera); Dragonlens (for dragonflies and damselflies, covering the order Odonata), and, the latest addition, Floralens (covering the Plantae kingdom).
Fig 1: Detail of the FloraOn web application.
@@
The methodology used to derive all the Biolens models has two core traits: (a) the use of data from research-grade public repositories for dataset construction, and (b) the use of Google’s AutoML Vision (GAMLV) to derive the actual models. This methodology is briefly described in a short scientific outreach article (in Portuguese) [12], and, more thoroughly, in MSc theses [13, 14] (both covering different stages of the work on Floralens), and a BSc project report [15] (covering Lepilens). Our evaluation of Floralens shows that, with a carefully designed dataset, current off-the-shelf machine-learning cloud-based services output models whose performance rivals or exceeds that obtained with models provided by state-of-the-art citizen science projects. Besides the baseline computer vision model, we study the impact of using multiple images by specimen and the geographical distribution of species on the model’s accuracy. We also compare the Floralens model to those provided by the state-of-the-art platform Pl@ntnet.
The main contributions of this paper are, thus, as follows:
  • a dataset for the Portuguese Flora available from Zenodo [16];
@@
  • a high-accuracy CNN-based model for the Portuguese native flora publicly available via web and mobile applications;
@@
  • a thorough quantitative evaluation of the model and comparison with the state-of-the-art platform Pl@ntNet;
  • a complete set of Python notebooks used to perform the analysis presented in this paper, also available from Zenodo.
@@
The remainder of this paper is structured as follows. Section 2 describes the current state-of-the-art regarding automatic taxonomic identification based on deep learning. Section 3 describes the construction of the dataset used in this study, the model’s generation using GAMLV, and the evaluation metrics used to gauge its accuracy. Section 4 describes the results obtained with the model and its refinements, such as using multiple images of the same specimen for classification and ecological indicators such as species’ geographical distribution. A detailed performance comparison with Pl@ntNet is also presented. Section 5 describes the software artifacts and datasets produced in the scope of this work. Finally, Section 6 discusses the main findings of this study and puts forward future research goals.
2 State-of-the-Art
@@
The advent of deep learning allowed the development of the first tools for automatically identifying plant species from input images [17, 18, 19]. Their continuous refinements quickly reached a point where automatic identification rivaled those made by specialists [20], thus attesting to the transformative role of AI in this field [21]. Nowadays, models based on deep learning are central tools in major citizen science platforms such as iNaturalist [2], Observation.org [3], and Pl@ntNet [6] and are integrated into web and mobile applications. Most of these models are derived using Convolutional Neural Networks [22, 23] and, recently, Vision Transformers [24, 25, 26]. Citizen science platforms generate important artifacts such as curated datasets, often made available to the public through biodiversity data portals, notably GBIF [11]. These datasets enable the development of other ML models, as is the case of Floralens that, in addition to FloraOn, uses data sampled from GBIF datasets provided by iNaturalist, Observation.org, and Pl@ntNet (cf. Section 3). However, the dataset construction and model derivation methodologies used by these citizen science platforms are not documented in detail in the literature: [2] makes reference to machine learning from an end-user perspective, some details can be found online in scattered form [27]; [6] provides a basic summary of the Pl@ntNet methodology to derive models, the information is also given in slightly more extended form in [28]; [3] is a short abstract concerning their mobile app ObsIdentify [29], and both the app and the site interface are enabled by an API [21, 30] whose underlying model is not explained. Moreover, extended access to the classification services in these platforms for research is hampered by the fact that in some cases the service is not exposed by the API or, commonly, requires the payment of a significant fee or subscription for large numbers of requests.
@@
Deriving a CNN-based model or other ML models often requires expert knowledge, non-trivial configuration aspects, writing specialized code, and a heavyweight computational infrastructure for storing data and model training. The strain is more acute when the amount of data involved is non-negligible, as in the case of Floralens where we use approximately 300,000 images to derive a model. AutoML cloud services address these needs, of which GAMLV is an example. Like Google, other major cloud providers have MLaaS offerings, e.g., Amazon Rekognition [31], Apple Create ML [32], and Azure AutoML [33]. These services allow the automation of several aspects of model derivation backed by computational infrastructures including special-purpose hardware such as GPUs or TPUs, necessary to derive a model in a reasonable amount of time. Models derived using these tools are referenced in the literature of several areas of knowledge, e.g., biology [34], medicine [35, 36], agriculture [37], engineering [38]. Comparative results between GAMLV and one or more competitor platforms are provided in some of these works [34, 35, 36], generally illustrating good performance by GAMLV-derived models. In this work, we used GAMLV since we had free access to the toolkit and computational infrastructure via Google Cloud’s Research Credits program.
This is a rapidly evolving research area stoked by developments in machine learning and available computational power and infrastructures. A recent trend is using metadata associated with observations to improve the identification precision. Pl@ntNet, for example, developed regional models by dividing the globe into several biogeographic domains [39] following published regional floras such as the WCVP/Kew [40] and introduced metadata providing information on the anatomic part of the plant depicted in an image, e.g., flower, leaf, stem. A similar approach is proposed by Rzanny et al. [41], where datasets with images taken from 5 different predefined perspectives per specimen are used. The authors then derive models for each perspective and fuse their outputs by averaging them across species to improve global precision. Mäder et al. [5] (Flora-Incognita) use a CNN to produce a model that classifies images of German native plants. A model that outputs the odds of finding a species in a given spatial and time context is also produced using a deep feedforward network. The outputs of these models are then fused using a recurrent neural network. More recently, Brun et al. [42] (FlorID) use data-efficient image transformers trained with images of specimens with associated locations, timestamps, and plant parts (e.g., vegetative, flower, fruit, trunk). As in Flora-Incognita, they use location (including elevation) and time attributes (season) to produce ecological predictor models that output the species most likely to be observed in a given spatial and temporal setting. They fuse the outputs of these ecological predictors with the image-based models to improve global precision. Despite these developments, some problems remain such as the apparent difficulty that these models have in identifying endangered species [43]. Floralens is closer in philosophy to regional flora projects such as Flora-Incognita [5] and FlorID [42].
Another trend that has generated considerable interest is the generation of global models based on extreme datasets such as iNat Challenge [44] and PlantCLEF/LifeCLEF [45, 46], consisting of millions of images representing tens of thousands of species out of the world’s $\sim$300K.
3 Methods
In this section we describe the creation of the Floralens dataset, the derivation of the Floralens model, and the evaluation metrics used to assess the latter’s performance.
3.1 Data
@@
The FloraOn repository contains geo-referenced records of Flora species with associated images. The image data is relatively broad in scope, as it covers 78% of the entire catalog (2,127 out of 2,712 species), but has a limited volume: on average there are just 11 images per species, and, unsurprisingly, our 50-image lower bound threshold is not met for a single species. Hence, to adequately populate the Floralens dataset, we retrieved the FloraOn images but also considered image data from three publicly-available datasets stored and made available at GBIF by three citizen science platforms: iNaturalist [47], Observation.org [48], and Pl@ntNet [49]. This approach is used by other platforms when building their datasets, e.g., FlorID [42].
The GBIF datasets provide validated observation data and associated images, originally submitted by users of the respective platforms. However, the validation process differs among data sources, as discussed further in this section. For each dataset and each species in our universe, we used GBIF portal queries to obtain observation records in the Darwin Core Archive format [50]. Each record corresponds to an observation of a specimen, typically made in the wild by citizen scientists or experts, accompanied by its taxonomic identification, date, geographical location, and one or more images. The GBIF portal queries were parameterized to cover continental Europe, including Great Britain, Ireland, and the Mediterranean islands. We did this as a preliminary analysis suggested that observations from Portugal or even the Iberian Peninsula would yield limited data in terms of volume and variety. Fortunately, many species of the Portuguese flora are widely distributed.
@@
Fig 2: Dataset histograms ($x$-axis: number of images; $y$-axis: number of species).
This source-based prioritization intends to define a dataset where images are less prone to identification errors. It takes into account the curation processes associated with each data source. FloraOn is curated by botanists and features high-quality images. These often feature subtle details that help secure the identification of a species. Pl@ntNet data goes through a curation process that involves machine learning, contributors’ reputation scores, and geo-based species verification [49]. Observation.org data can result from automatic validation through image recognition coupled with a check for other approved observations in the geographical vicinity, or through an expert volunteer when automated validation fails [48, 51]. Finally, iNaturalist identifications result from a crowd-sourcing effort whereby a “research-grade” identification for an image can be obtained from the consensus of just two citizen scientists [47, 52].
We used images without distinction of plant features as this metadata was not available from the curated datasets. Similarly, we do not attempt to gauge “image quality”. The citizen-science platforms provide guidelines for submitting acceptable quality images, and their curation processes mostly filter inadequate items (e.g., out of focus, under/over-exposure, presence of significant digital noise, insufficient resolution, subject not sufficiently separated from surrounding vegetation).
@@ 3.2 Model
Fig 3: Model derivation using GAMLV.
@@
Fig 4: GAMLV interface for model training and deployment.
@@
Fig 5: Layers of the CNN model (fragment).
@@
The CNN architectures at stake are picked from the MnasNet family [58], developed with mobile and embedded devices in mind. The high-level choice between models offered by GAMLV (c.f. Figure 4(a)) corresponds to three different MnasNet instantiations that do not differ in structure, just in the density of connections between layers and the number of internal weights.
3.3 Evaluation Metrics
The performance of the Floralens model was evaluated over several test datasets of observations using the following metrics. Precision is the ratio of true positives (TP) relative to the total number of positives (TP $+$ FP). A positive (identification) occurs when the classification score (a probability) returned by the model equals or exceeds a confidence level set as the threshold for analysis. Recall is the ratio of true positives relative to the total number of true examples (TP $+$ FN). The confidence level is relevant for practical uses of a model. Using a low confidence level means that false positives are tolerable for the intended use, and we will get lower precision. On the other hand, a value too high for the confidence level will result in more false negatives, hence lower recall.
@@ 3.3 Evaluation Metrics
4 Results
In this section, we present the results of multiple experiments conducted to characterize the Floralens model’s performance and compare it to state-of-the-art models. In the following subsections, larger values indicate better model performance.
4.1 Base Results
Figure 6 shows the results for precision and recall for the Floralens model applied to the test set given in Table 2, more precisely the macro-average of precision and recall values for all species. The area-under-curve (AUC) for the precision-recall correlation (in 6(a)), also known as the average precision, is 0.72 (the maximum value would be 1.0). Putting the confidence levels in perspective (in 6(b)) we see that precision and recall are approximately equal to 0.7 for a confidence level of 0.2. For a confidence level of 0.5 precision equals 0.85 and recall equals 0.53. Overall, the results indicate a reasonable predictive power for the Floralens model.
@@ 4.1 Base Results
Fig 6: FLTS results: precision and recall.
(a) Overall results and results per data source.
(b) Distribution of results per species.
Fig 7: FLTS results: Top-1, Top-5 and MRR
(a) Results according to species growth form.
(b) Results for special categories.
Fig 8: Results according to species growth form and special categories.
Figure 7 shows the Top-1, Top-5, and MRR results for FLTS in two plots: (7(a)) overall results, and per data source used in constructing the dataset (FloraOn, iNaturalist, Observation.org, and Pl@ntNet) and; (7(b)) boxplots displaying the distribution of the metric results over the universe of species, i.e., values for the metrics at stake for each species. The overall results again indicate reasonably good predictive power: 0.67 for Top-1 (i.e., roughly two-thirds of the FLTS images are correctly classified with rank 1), 0.86 for Top-5, and 0.75 for MRR; the results also indicate a relatively homogeneous predictive power across all data sources and metrics, as the maximum difference in values between them does not exceed 0.06: a Top-1 value of 0.64 for Observation.org vs. corresponding values of 0.70 for FloraOn and iNaturalist. Regarding the distribution of results per species, the median values for Top-1, Top-5, and MRR are respectively 0.70, 0.89, and 0.75. Looking at the first quartile, we also get indicators for reasonable predictive power: only 25 % (resp. 75 %) of species have a Top-1 value of 0.55 or lower (resp. or higher), and; the first quartile values for Top-5 and MRR are 0.80 and 0.65.
In Figure 8 we present two types of complementary results. The plots detail the Top-1, Top-5 and MRR grouped by: (8(a)) species growth type and; (8(b)) the special categories of species that are endangered (part of the IUCN red list), endemic, protected, or rare. In both cases, the FloraOn site [10] is the reference for the presented subsets of species. The number of species in each subset is given between parenthesis in the figures (e.g., 42 species of ferns). The y-axis reference ticks correspond to the overall FLTS values for the Top-1 (0.67), MRR (0.75), and Top-5 (0.86) metrics. Regarding species growth type, we can observe that grass species have the worst performance, as they tend to be hard to distinguish visually, and tree species are also slightly below the overall FTLS performance. As for the special categories, none of them stand out below or above the overall FTLS values. This is a good trait for the endemic species since these can be challenging for models to detect [43].